A transformer consists of a 500-tum primary coil a
Answer #1

Here ,

for the transformer

let the secondary currnet is Is

Ip = 3 A

Np = 500

Ns = 2000

as for an ideal transformer

Is * Ns = Np * Ip

3 * 500 = 2000 * Is

Is = 0.750 A

the current in the secondary is a) 0.075 A